# Pulse-PPG: An Open-Source Field-Trained PPG Foundation Model for Wearable Applications Across Lab and Field Settings

### v1

@@ 5.3.3. Results
Figure 4 visualizes the difference in performance between Pulse-PPG pre-trained on the wearable field PPGs vs. pre-trained on clinical PPGs. Full results, including comparisons of PaPaGei pre-training with wearable field vs. clinical data can be found in Table 3 in the Appendix, while referencing the second row named ”Pre-Train Data”.
Pulse-PPG consistently exhibits stronger performance on 10 out of 11 tasks when pre-trained on the Field Wearable PPG data versus when pre-trained on the Clean Clinical PPG. For all other tasks except for sleep disturbance, the field wearable pre-training yielded consistently stronger results, sometimes substantially so. For example, for Field Wearable datasets, F1 score for stress improved by .0848 and F1 score for activity improved by .1218. Even for the clinical tasks, we see that pre-training on the wearable PPGs achieves much stronger performance (i.e., .10 F1 improvement in hypertension prediction and $\sim 1$ MAE improvement in systolic and diastolic BP regression). In some cases, such as in the WESAD evaluation datasets, one metric was worse, but all of the remaining metrics were stronger.
Pre-training on field wearable data appears to yield a more robust and generalizable representation, outperforming clinical pre-training—even on clinical tasks. Although performance gains in clinical tasks are modest, the consistency demonstrates that pre-training on field wearable data performs as well as, and sometimes better than, pre-training on clinical data. As discussed earlier in Section 5.2.3, relatively poor results on sleep can be explained by lack of sleep data present in the MOODS dataset. This trend is also observed in PaPaGei, though to a lesser extent. When pre-trained on wearable data, PaPaGei outperforms clinical pre-training in 6 out of 11 tasks, including five wearable PPG evaluation datasets (Activity - MOODS, Instant HR - PPG-DaLiA, Activity - PPG-DaLiA, Stress (2) - WESAD, and Stress (4) - WESAD). This aligns with the idea that pre-training and evaluating on wearable data introduces less domain shift. The difference in performances across pre-training settings between Pulse-PPG and PaPaGei can be attributed to their pre-training strategies. PaPaGei relies on PPG-specific morphologies, which may not generalize well across domains with varying PPG patterns (e.g. wearable vs. clinical). In contrast, Pulse-PPG leverages a learnable motif-similarity distance, enabling it to capture more general PPG patterns, which are applicable across different settings.
5.4. Experiment 3: Evaluation of Pulse-PPG Use-cases
@@ 5.4.1. Background
In this section we quantitatively measure the effectiveness of Pulse-PPG for 11 downstream tasks, to provide further insight into its utility to the research community.
@@ 5.4.2. Experimental Design
  • Linear Probe: The model’s embeddings are frozen, and a logistic classifier or linear regression is trained on the embedding vectors. This is a simple evaluation designed to evaluate the quality of the learned representations to generalize to new tasks. By requiring minimal computational resources, it also enables a broader range of researchers to utilize the model’s learned representations.
  • Fine Tune: This evaluation is designed to boost task-specific performance by training the entire network end-to-end to adapt to a specific task. This highlights the adaptability of the foundation model, showing its potential to serve as a strong starting point for specialized tasks.
In addition to this, we include a naive baseline for clear comparison. The naive classifier always predicts the majority class, and the naive regressor predicts the training mean. Though simple, this benchmark contextualizes our model’s performance, highlighting the value of learned representations and ensuring transparent evaluation.
Table 2. Performance metrics for the Pulse-PPG foundation model across diverse wearable and clinical datasets. There are three evaluation approaches: (1) Naive methods, which serve as a baseline; (2) Linear Probing, where the model’s embeddings are frozen and a simple classifier or regressor is trained, enabling a fair assessment of the general-purpose representations; and (3) Fine-Tuning, which adapts the entire network end-to-end for improved task-specific performance.
PPG RelCon
Pre-Train Data: - Wearable Field
Eval Method: Naive Linear Probe Fine-Tune
@@
Wearable Field Stress (2) MOODS F1 Score 0.3859 0.5398 0.5872
@@
Clinical Sleep Disturbance (2) SDB F1 Score 0.3945 0.4630 0.5782
@@
Average HR (R) PPG-BP MAE 7.965 4.003 3.697
MSE 98.08 27.02 23.26
MAPE 0.1176 0.0573 0.0531
Hypertension (2) PPG-BP F1 Score 0.4459 0.6975 0.7213
Accuracy 0.8049 0.8130 0.7967
Precision 0.4024 0.7009 0.7031
Recall 0.5000 0.6944 0.7633
AUPRC 0.1951 0.4266 0.7408
AUROC 0.5000 0.7971 0.8258
5.4.3. Results
Table 2 shows the full set of results.
Overall, the results demonstrate that Pulse-PPG consistently outperforms the naive baseline across all tasks. When evaluated with a linear probe, the model’s frozen embeddings already yield competitive performance, highlighting the strength of the learned representations. For example, in the wearable field stress detection task, a traditionally difficult task, especially in the field setting, where noise and variability are inherent, a simple linear probe is able to achieve an F1 score of 0.5398, compared to the naive method achieving an F1 of 0.3859. Additionally, in the binary stress classification for the WESAD dataset, the simple logistic regression on our learned feature can achieve an F1 of 0.8759, more than doubling the naive model’s F1 score of 0.4065.
Fine-tuning further enhances performance by adapting the entire network to the specific task. This is evident across all tasks, where fine-tuning yields the best results. Although occasionally, linear probing may outperform fine-tuning for one metric, for the remaining metrics, the fine-tuned results are stronger. This underscores the model’s adaptability and its potential to serve as a robust starting point for researchers to adapt for their specialized applications.
The linear probe evaluation offers a computationally efficient means to leverage powerful representations, while fine-tuning maximizes task-specific performance. Together, they highlight the versatility and practical utility of foundation models in both exploratory and application-focused research settings.
6. Discussions
6.1. Shifting the Paradigm: From Task Specific Models to Foundation Models
Early mHealth research was limited by the lack of large-scale field datasets, prompting researchers to collect data in controlled lab environments. A common approach involved recruiting a group of participants, collecting physiological data in a lab setting, training a model on this clean data with clean labels, and then deploying the model on field data from the same participants (Ertin et al., 2011; Plarre et al., 2011). This systematic approach minimizes variability between training and deployment and generally results in better performance as shown in recent works (Toshnazarov et al., 2024). For greater between-person generalizability, researchers trained the model on one group of lab participants and tested on a separate group of lap participants (Hovsepian et al., 2015). For lab-to-field generalizability, the model was tested on a third group of participants who collected data in field settings (Hovsepian et al., 2015). A further step toward real-world applicability involved training AI models directly on growing volumes of field data and testing them on the same participants using 10-fold cross-validation or cross-subject validation (Mishra et al., 2018; Zhang et al., 2024a). This approach posed significant challenges due to the inherent noise and variability in real-world data, making model development more complex. The majority of these works (Mishra et al., 2018; Zhang et al., 2024a) report F1 scores of around 0.5 for stress classification when using PPG only data.
Further, a key limitation of task-specific models is their inflexibility—adapting to even a slightly different task (e.g., transitioning from binary to multi-class classification) often requires substantial retraining and modifications to the modeling pipeline. To overcome these challenges, a more recent paradigm has emerged: foundation models that are task-agnostic and general-purpose. These models leverage large-scale datasets to learn generalizable representations, enabling adaptation to diverse downstream tasks without requiring task-specific redesigns. Our results show that this approach has a high potential to improve performance on a wide variety of mHealth tasks in the field environment.
6.2. Field-to-Lab Generalizability
Cleaner data has been a major advantage of lab data. For detecting psychological states such as stress, another major advantage of lab data is clean labels. In the field conditions, the labels are usually self-reported. They do not temporally align with the physiological data due to a large time lag between how quickly physiology recovers versus the perception of stress. Further, unlike lab data where an entire block of data corresponds to a stress or non-stress task, blocks of data receiving the same label can consist of a mixture of stress and non-stress segments, due to lack of control over the experiences of a participant. These label impurities make it difficult to train task-specific models that use the labels together with the data.
Unlike traditional task-specific models (Choi et al., 2022; Hasanpoor et al., 2022; Motaman et al., 2022; Alshareef et al., 2022; Mitro et al., 2023), self-supervised learning-based models are designed to reduce or eliminate the dependency on such explicit labels (Spathis et al., 2022; Haresamudram et al., 2022; Rabbani and Khan, 2022). This is achieved through their inherent structure, where the pre-training process derives labels directly from the data itself, without needing human-provided annotations. This self-supervised approach represents a significant departure from conventional methods, as it taps into the underlying patterns of the data rather than relying on predefined labels. This shift in model design has paved the way for an exploration of field-to-lab generalizability. Results from our pre-trained foundation model on field data suggest that this approach not only enhances performance on field-based mobile health tasks but also demonstrates unexpected success in clinical and laboratory settings, which are typically more controlled and structured.
To our knowledge, this is the first work that demonstrates the benefits of pre-training a foundation model on field data for improving performance not just in real-world, field-based mHealth tasks but also in more traditional clinical and lab environments. The success of the model across such diverse task domains—ranging from clinical and mHealth lab settings to field tasks involving PPG data—challenges the conventional belief that models must be trained on clean, highly curated datasets to perform well in clinical and lab settings. This finding opens up new avenues for leveraging field data to develop more robust, generalized models capable of performing across a wide range of environments.
6.3. Optimizing Task Specific Performance
Previous research primarily focused on task-specific performances. The foundation models, on the other hand, are general-purpose. But, an added benefit of foundation models is that they can also be optimized for enhancing the performance of specific tasks. When the focus is more on enhancing task-specific performance with pre-trained models, instead of using a linear probe approach used to test generalizability, which keeps the pre-trained model weights frozen and simply adds a linear layer on top for training, fine-tuning is the more appropriate approach that involves training the model end-to-end, using the pre-trained model weights as a starting point. This allows the model to adapt to the task-specific characteristics of the data more closely, which enables it to perform better for that task. Additionally, future work can explore integrating additional data modalities, such as accelerometer data, to more effectively identify when the PPG is affected explicitly by motion artifacts. This could enable our model to learn a representation that can better isolate physical activities and mental stressors. Beyond motion data, contextual features—such as activity level, time of day, or environmental conditions—can enable more comprehensive and informative feature extractions (Mishra et al., 2018; Toshnazarov et al., 2024). Finally, selectively incorporating demographic factors, while ensuring fairness and minimizing bias, may provide valuable insights into individual differences in physiological responses. Our work provides a new starting point, enabling the community to build more comprehensive multimodal representations for more accurate task-specific results.
7. Limitations
7.1. Stress Downstream Task in Field Conditions
One limitation of this study is that the stress episodes used for the field downstream task of stress vs. non-stress classification were obtained from a dataset (Neupane et al., 2024) that used a commercial pre-trained model to obtain self-reported stress ratings. Other datasets have employed different approaches to obtain stress ratings (Toshnazarov et al., 2024; Zhang et al., 2024a). Although the competing foundation model was also tested on the same dataset, testing the model performance on multiple large field datasets can demonstrate greater generalizability for stress detection in the field.
7.2. Expanding Pre-training Data for Broader Generalization
The wearable field dataset used to pre-train the foundation models in this study originated from a single field study involving a diverse and general population from throughout the U.S. (Neupane et al., 2024). While this approach demonstrated the model’s utility, relying on data from one study inherently limits the diversity of field conditions, stress episodes, and physiological responses. To be more inclusive and more robust to field conditions, future work should incorporate data from multiple studies across varied populations, geographic regions, and contexts, to enhance generalizability. Expanding the diversity of training data can improve the model’s adaptability to new environments, mitigate biases from a single study, and ensure more robust and equitable performance across broader applications.
8. Conclusion
Our work presents Pulse-PPG, the first open-source1 PPG foundation model pre-trained on field-collected data, underscoring the crucial role of real-world variability in enhancing model robustness and applicability. By outperforming state-of-the-art foundation models trained on wearable lab or clinical data across multiple downstream tasks, Pulse-PPG establishes the necessity of incorporating real-world variability into model training. Additionally, our findings indicate that field-trained models can perform competitively even in controlled lab datasets, challenging conventional assumptions about the superiority of lab-trained models. The open-source1 release of Pulse-PPG will allow for further advancements in PPG-based health monitoring, enabling the research community to build upon a model that is inherently designed for real-world deployment. This work highlights the importance of shifting towards field-data-driven AI models to bridge the gap between controlled research settings and practical healthcare applications. Future research should explore ways to refine field-trained models, investigate hybrid training approaches, and expand the applicability of foundation models across diverse physiological signals.
Appendix A Appendix
A.1. Evaluation Dataset Details
Cardiovascular health is a key area of focus, with biomarkers such as systolic blood pressure, diastolic blood pressure, and heart rate serving as vital indicators. The PPG-BP dataset includes short PPG recordings from 219 participants to acquire information on basic physiology of individuals. As part of the data collection process participants were asked to relax for about 10 minutes. After that, both blood pressure and finger PPG was collected within the next 3 minutes with 3 short PPG recording of 2.1 seconds each per participant. , most of whom have hypertension, along with corresponding blood pressure and heart rate measurements (Liang et al., 2018). Using this dataset, we estimate these biomarkers and predict hypertension.
Stress and affective states, which significantly impact physical and mental health, are evaluated using the WESAD dataset (Schmidt et al., 2018b). The WESAD dataset collected physiological signals from 15 participants (12 males, 3 females) using RespiBAN Professional (biosignalsplux, 2019) on the chest and an Empatica E4 (McCarthy et al., 2016) on the wrist. The dataset includes ECG, Blood Volume Pulse (BVP), Electrodermal Activity (EDA), and Skin Temperature (TEMP) collected at 700 Hz, 64 Hz, 4 Hz, and 4 Hz, respectively. The study protocol had five sessions: Baseline, Stress, Amusement, Meditation and Rest. In the baseline session, participants engaging in neutral activities for 20 minutes. Stress was induced using the Trier Social Stress for about 10 minutes. During the session, participants had to give a public speech and complete a mental arithmetic task for 5 minutes each. For the amusement session, participants watched funny video clips for 6.5 minutes. To facilitate recovery, a meditation session followed both the stress and amusement sessions, along with an additional 10-minute rest period after the stress session. As part of the downstream evaluation tasks, we perform both binary and multi-class classification using only PPG data collected during the sessions. For binary classification, we combined baseline and amusement sessions to form non-stress class while stress class was the original stress session. For multi-class classification, we use Stress, Baseline, Amusement and Mediation sessions. For both types of classifications, each session was divided into non-overlapping 1-minute windows. Hence, stress vs. non-stress class had approximately 10 vs. 26 windows for the binary classification, while for the multi-class classification, window distribution was 20, 10, 6, and 7 for Baseline, Stress, Amusement and Mediation respectively.
Physical activity, which is a key indicator of one’s healthy lifestyle, is assessed using the PPG-DaLiA dataset (Reiss et al., 2019b). This publicly available multimodal dataset comprises physiological and motion data collected from wrist- and chest-worn wearable devices. It includes recordings from 15 participants (seven males and eight females) aged 21 to 55 years, performing eight distinct activities (sitting still, ascending/descending stairs, playing table soccer, cycling, driving, taking lunch breaks, walking, and working). Additionally, transition periods between activities were labeled as a separate ‘zero’ activity. All activities were conducted in conditions closely resembling real-life scenarios, with each participant’s recorded session lasting approximately 2.5 hours. Data were collected using the RespiBAN Professional and Empatica E4 wearable devices. The dataset includes PPG recordings and heart rate measurements during various activities. We utilize this dataset for two key tasks: activity classification (9 classes) and instantaneous heart rate estimation.
Sleep-disordered breathing (SDB) can lead to significant health complications in both adults and children. In children, SDB is associated with issues ranging from daytime sleepiness to severe conditions like developmental delays and growth failure. To examine its impact, the SDB dataset (Garde et al., 2014) collected physiological data from 146 children using a Phone Oximeter (Karlen et al., 2011), which recorded PPG at 62.5 Hz and SpO2 at 1 Hz during six hours of overnight sleep. Participants were categorized into SDB-positive and SDB-negative groups based on their Apnea-Hypopnea Index (AHI) scores, which quantify the severity of obstructive sleep apnea. Following the approach in (Pillai et al., 2024), we utilize these AHI ratings to frame SDB prediction as a binary classification task.
Field Dataset: Stress Detection and Motion Detection In the field study, participants received 5.2 prompts on average per day for annotation from which they responded to 3.9, yielding a response ratio of 75% using a Likert scale with response items (‘Not stressed,’ ‘Probably not stressed,’ ‘Unsure,’ ‘Probably Stressed,’ ‘Stressed’). If the annotation was either ‘Stressed,’ ‘Probably Stressed,’ or ‘Unsure’, the app prompted them to provide the semantic location and the likely stressor. From a total of 128,006 physiological events detected by an AI model running on the smartwatch, 35,785 events were presented to participants for annotation, of which participants annotated 26,682 events. Following prior research (Toshnazarov et al., 2024; Mishra et al., 2018), we use 1-minute windows for the downstream stress vs. non-stress detection task in the field. Additionally, to address label noise, we remove participants who exhibit no variation in their stress ratings or those with a low compliance rate in submitting stress reports. Specifically, we exclude three participants due to no variation in their ratings and 16 participants with low compliance compared to the rest. This filtering results in a revised train/validation/test split of 74/14/13 participants. For classification, we group ‘Stressed,’ ‘Probably Stressed,’‘Unsure’ into the Stress class, as participants provided stressor information for these ratings. Meanwhile, ‘Not stressed,’ ‘Probably not stressed’ are categorized under Non-Stress. The final distribution of Stress vs. Non-Stress classes is approximately 58% and 42%, respectively, across a total of 206,088 one-minute labeled stress and non-stress windows.
For the other field task we chose stationary vs. motion as the downstream task. To do so, we use the first 10 days of each participant’s data, assuming this duration provides enough representation of an individual’s stationary and motion behavior. Since the field dataset lacks explicit activity labels, we employ a pre-trained activity recognition model (Saleheen et al., 2021) to classify 20-second segments of accelerometry data into five activity categories: ‘Stationary,’ ‘Walking’, ‘Stairs’, ‘Sports’, and ‘Exercise’. We then group ‘Walking,’ ‘Stairs,’ ‘Sports,’ and ‘Exercise’ into the Non-stationary class. From these 10 days of data, we obtain a total of 3,559,429 stationary and non-stationary windows, with a class distribution of 15% stationary and 85% non-stationary. We extract raw PPG data corresponding to these labeled 20-second intervals for each participant. To maintain consistency with prior experiments, we use the same original split of 84/18/18 to partition the total 3,559,429 stationary and non-stationary windows for the downstream task.
This benchmarking approach will demonstrate the versatility of our PPG foundation model, showcasing its ability to adapt to the domain shift of various datasets, and generalize well across diverse health-related tasks, notably after learning from noisy PPG data from a field dataset.
A.2. Evaluation Methodology Details
All PPG signals were re-sampled to match the 50 Hz sampling frequency present in our pre-training MOODS dataset (Neupane et al., 2024). For our evaluations using PaPaGei, we made sure to resample the signals to 125 Hz in order to match the sampling frequency that they trained on (Pillai et al., 2024).
In order to train and evaluate our linear probes, we combined the train and validation splits and used a cross validation hyperparameter grid search with the Scikit Learn package. For the regression tasks, we utilized a linear regression with an L2 regularization and conducted a grid search over the following parameters: ’alpha’:[0.1,1.0,10.0,100.0],’solver’:[’auto’,’cholesky’,’sparse_cg’] with the scoring function as macro F1. For the classification tasks, we utilized a logistic regression with an L2 regularization and lbfgs solver and conducted a grid search over the following parameters ’C’:[0.01,0.1,1,10,100],’solver’:[’lbfgs’],’max_iter’:[1000,10_000] with the scoring function as negative MSE. Inputs into the probes were normalized with the Scikit Learn StandardScaler module.
In order to train and evaluate our fine-tuning results, we fit a classification or regression head ontop of the existing network, then train the entire network end-to-end. For classification, the head was made up of a layer normalization layer, followed by a linear layer, with the loss function being a linear combination of Cross Entropy loss, a soft F1 loss (Bénédict et al., 2021), and Dice loss (Li et al., 2019). For regression, the head was an MLP layer, with an initial layer with an output dimensionality of 128, followed by the GeLU activation, followed by a final linear layer. The loss function used was the MSE loss function. For both evaluations, we utilize an Adam optimizer. The checkpoint with the highest macro F1 score and the lowest MAE on the validation split were selected for inference on the test set.
For more specific implementation details, please check our public1 codebase.
A.3. Additional Results
The full results for comparing Pulse-PPG vs. PaPaGei and comparing Pulse-PPG pre-trained on Wearable Field vs. Clinical and comparing PaPaGei pre-trained on Wearable Field vs. Clinical can be found in Table 3 below.
Table 3. Comparing Pulse-PPG vs. PaPaGei and Wearable Field PPG vs. Clinical Clean PPG pre-training data.
PPG RelCon PaPaGei
Pre-Train Data: Wearable Field Clinical Wearable Field Clinical
Eval Method: Linear Probe
Wearable Lab Instant HR (R) PPG-DaLiA MAE 8.936 18.60 11.37 19.18
MSE 149.3 498.2 212.9 525.7
MAPE 0.1166 0.2581 0.1522 0.2708
Activity (9) PPG-DaLiA F1 Score 0.3039 0.2250 0.2884 0.1861
Accuracy 0.4104 0.3276 0.3862 0.3020
Precision 0.3882 0.2631 0.3519 0.2142
Recall 0.3107 0.2457 0.3016 0.2041
AUPRC 0.3504 0.2527 0.3313 0.2149
AUROC 0.8051 0.7741 0.7964 0.7438
Stress (2) WESAD F1 Score 0.8759 0.8009 0.8237 0.6896
Accuracy 0.8904 0.8493 0.8356 0.7260
Precision 0.8688 0.8795 0.8149 0.6863
Recall 0.8848 0.7726 0.8565 0.6943
AUPRC 0.9410 0.8724 0.7518 0.6465
AUROC 0.9687 0.9374 0.9130 0.8043
Stress (4) WESAD F1 Score 0.5431 0.5126 0.4976 0.4010
Accuracy 0.6517 0.6629 0.5393 0.5169
Precision 0.6003 0.5124 0.4976 0.3947
Recall 0.5716 0.5387 0.5258 0.4079
AUPRC 0.5671 0.6931 0.4813 0.5202
AUROC 0.7807 0.8456 0.7257 0.7684
Wearable Field Stress (2) MOODS F1 Score 0.5398 0.4550 0.4141 0.4398
Accuracy 0.6156 0.6135 0.6251 0.6257
Precision 0.5606 0.5244 0.5355 0.5538
Recall 0.5460 0.5085 0.5042 0.5114
AUPRC 0.4275 0.3960 0.3923 0.4030
AUROC 0.5691 0.5382 0.5275 0.5246
Activity (2) MOODS F1 Score 0.6859 0.5641 0.5375 0.5318
Accuracy 0.8705 0.8508 0.8483 0.8458
Precision 0.7835 0.7401 0.7349 0.7058
Recall 0.6509 0.5562 0.5401 0.5365
AUPRC 0.5835 0.3916 0.3673 0.3410
AUROC 0.8722 0.7205 0.7302 0.6941
Clinical Sleep Disturbance (2) SDB F1 Score 0.4630 0.5005 0.5012 0.5588
Accuracy 0.5364 0.5898 0.5808 0.6609
Precision 0.4633 0.5108 0.5075 0.6079
Recall 0.4673 0.5084 0.5062 0.5672
AUPRC 0.3085 0.3636 0.3577 0.4236
AUROC 0.4437 0.5122 0.5088 0.5528
Systolic BP (R) PPG-BP MAE 13.62 14.60 14.49 13.99
MSE 286.5 339.0 338.2 321.1
MAPE 0.1065 0.1138 0.1131 0.1096
Diastolic BP (R) PPG-BP MAE 8.878 9.415 9.848 9.246
MSE 118.2 136.7 135.3 127.0
MAPE 0.1232 0.1299 0.1374 0.1290
Average HR (R) PPG-BP MAE 4.003 4.258 3.971 4.549
MSE 27.02 27.60 25.74 32.59
MAPE 0.0573 0.0607 0.0575 0.0662
Hypertension (2) PPG-BP F1 Score 0.6975 0.5939 0.5839 0.6403
Accuracy 0.8130 0.7073 0.6829 0.7317
Precision 0.7009 0.5887 0.5824 0.6311
Recall 0.6944 0.6130 0.6136 0.6755
AUPRC 0.4266 0.3051 0.3413 0.3214
AUROC 0.7971 0.6936 0.7100 0.7201

### v2

@@ 6.3.3. Results
Figure 4 visualizes the difference in performance between Pulse-PPG pre-trained on the wearable field vs. clinical PPGs. Full results, including comparisons of PaPaGei pre-training with wearable field vs. clinical data, can be found in Table 7 in the Appendix, while referencing the second row named ”Pre-Train Data”.
Pulse-PPG consistently exhibits stronger performance on 10 out of 11 tasks when pre-trained on the Field Wearable PPG data versus when pre-trained on the clean Clinical PPG. For all other tasks except for sleep disturbance, the field wearable pre-training yielded consistently stronger results, sometimes substantially so. For example, for Field Wearable datasets, the F1 score for stress improved by .0848, and the F1 score for activity improved by .1218. Even for the clinical tasks, we see that pre-training on the wearable PPGs achieves much stronger performance (i.e., .10 F1 improvement in hypertension prediction and $\sim 1$ MAE improvement in systolic and diastolic BP regression). In some cases, such as in the WESAD evaluation datasets, one metric was worse, but all of the remaining metrics were stronger.
Pre-training Pulse-PPG on field wearable data appears to yield a more robust and generalizable representation, outperforming clinical pre-training, even on clinical tasks. Although performance gains in clinical tasks are modest, the consistency demonstrates that pre-training on field wearable data performs as well as, and sometimes better than, pre-training on clinical data. As discussed earlier in Section 6.1.3, relatively poor results on sleep can be explained by the lack of sleep data present in the MOODS dataset.
This trend of stronger performance when trained on wearable data is also observed in PaPaGei, though to a lesser extent. When pre-trained on wearable data, PaPaGei outperforms clinical pre-training in 6 out of 11 tasks, including five wearable PPG evaluation datasets (Activity - MOODS, Instant HR - PPG-DaLiA, Activity - PPG-DaLiA, Stress (2) - WESAD, and Stress (4) - WESAD). This aligns with the idea that pre-training and evaluating on wearable data introduces less domain shift. The difference in performance across pre-training settings between Pulse-PPG and PaPaGei can be attributed to their pre-training strategies. PaPaGei relies on PPG-specific morphologies, which may not generalize well across domains with varying PPG patterns (e.g., wearable vs. clinical). In contrast, Pulse-PPG leverages a learnable motif-similarity distance, enabling it to capture more general PPG patterns, which are applicable across different settings.
6.4. Experiment 4: Evaluation of Pulse-PPG Use-cases
@@ 6.4.1. Background
In this section, we quantitatively measure the effectiveness of Pulse-PPG for 11 downstream tasks, to provide further insight into its utility to the research community.
@@ 6.4.2. Experimental Design
  • Linear Probe: The model’s embeddings are frozen, and a logistic classifier or linear regression is trained on the embedding vectors, consistent with prior foundation model work (Abbaspourazad et al., 2023; Devlin et al., 2018; Oquab et al., 2023). This is a simple evaluation designed to evaluate the quality of the learned representations to generalize to new tasks. By requiring minimal computational resources, it also enables a broader range of researchers to utilize the model’s learned representations. Introducing a more sophisticated traditional-ML-based probe would confound the results, with it being more difficult to isolate representation quality from the strength of the classifier.
  • Fine Tune: This evaluation is designed to boost task-specific performance by training the entire network end-to-end to adapt to a specific task. This highlights the adaptability of the foundation model, showing its potential to serve as a strong starting point for specialized tasks. While computationally intensive, it is essential for applications requiring peak performance.
Table 3. Pulse-PPG performance across wearable and clinical PPG tasks. There are three evals: Naive, predicting majority class or mean; Linear Probing, which trains a simple head on the Pulse-PPG’s frozen embeddings; and Fine-Tuning, which trains Pulse-PPG’s network end-to-end.
Pulse-PPG
Eval Method: Naive LP FT
@@
Wearable Field Stress (2) MOODS F1 0.3859 0.5398 0.5872
@@
Clinical Sleep Disturbance (2) SDB F1 0.3945 0.4630 0.5782
@@
Average HR (R) PPG-BP MAE 7.965